There was not a section on combining like terms. I do,however know that inverse operations plays a big role in combining like terms. My question is how to you combine like terms?
Answer by lauramg012(4)

An example of combining like terms: x+x=2x
x could equal 10. It is the same as saying 10+10=2(10).
An example of unlike terms are x and y. You cannot combine x and y unless the values for x and y are given.
For example, if x=2y and y=10, then x+y = 2y+10 by the substitution property of equality.
y=10, so 2y+10 = 2(10)+10
20+10=30
However, if you were not given x and y values, then you could not combine the terms because they are not alike. x+y cannot equal 2xy.